There are three cell phone models in a store. When selecting a new cell phone, 25% of the customers choose model A, 33% choose model B, and 32% choose model C. The remaining customers buy from an old collection on which the average profit is $50. If the average profit earned on models A, B, and C is $60, $75, and $40, respectively, what is the expected value of the profit earned on all models?
1
Expert's answer
2019-04-17T12:53:13-0400


 

The expected value of the profit:



"E(x)=\sum(x_i*p_i)"

where xi is the value of the profit, pi is the probability of the value of the profit.




"E(x)=60*0.25+75*0.33+40*0.32+50*0.1=$57.55"